OpenStudy (anonymous):

Solve r^2+4=0. I totally forgot how to do this. HAHA.

OpenStudy (anonymous):

there is no real solution, because you have the sum of two squares there are two complex solutions though

OpenStudy (anonymous):

\[r^2+4=0\\r^2=-4\\r=\pm\sqrt{-4}=\pm2i\]

OpenStudy (anonymous):

This is what I looked for. Thanks.